# Active Learning for Control-Oriented Identification of Nonlinear Systems

### v1

@@ A.2 Improved Rates via the Delta Method
$\tau_{\mathsf{cov\,conc}}(\delta)=\frac{52T^{2}L_{f}^{4}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)(3d_{\phi}+\log(1/\delta))}{\lambda_{\min}(\Sigma^{\pi})^{2}}.$
If $K\geq\tau_{\mathsf{cov\,conc}}(\delta)$, we may invoke Lemma A.9 to show that with probability at least $41-\delta/$ 4,
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}),$
where $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}\leq\frac{1}{2}$. As a result, we have that $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$. Given this condition and the burn-in conditions $K\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/16),\tau_{\mathsf{small\,error}}(\delta/4),\tau_{\mathsf{h.o.t.}}(\delta/4)\right\}\mathclose{}$, we satisfy the conditions of Lemma A.5. Therefore, conditioned on the event of covariance concentration, we have that with probability at least $41-\delta/$ 4,
@@
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}$. Then by a triangle inequality and submultiplicativity, we have that under the success events of covariance concentration and the delta method,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 2\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}+2\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}$ (31)
In particular, we have reduced the parameter recovery errort to the sum of weighted norms of a self-normalized martingale. Calling on self-normalized martingale machinary Abbasi-Yadkori et al. (2011), Wagenmaker et al. (2021) we may bound each of these norms with high probability. In particular, invoking Lemma A.7 with $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}$, we find that with probability at least $21-\delta/$ 2, the following bounds hold:
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
Substituting these bounds into (31), we find that by union bounding over the success events, we have that with probability at least $1-\delta$,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 40\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1}}{K}.$
@@ Appendix B Results for smooth nonlinear systems
In particular, we first define $p_{w}$ to denote the density of the noise, $W_{t}$. Let $p_{\phi,\theta}(\cdot)$ denote the density over the trajectories $(X_{1:T+1},U_{1:T})$ induced by playing controller $\pi^{\theta}$ on the system (1) with parameter $\phi$. Observe that
$p_{\theta,\phi}(X_{1:T+1},U_{1:T})=\prod_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi^{\theta}_{t}(X_{t});\phi^{\star})).$ (32)
@@
$J(\pi^{\theta},\phi)=\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
@@
$\frac{d}{dx_{1}}J(\pi^{\theta_{\bm{s}}},\phi_{\bm{m}})=\int\frac{d}{dx_{1}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
$=\int\underbrace{\left(\frac{d}{dx_{1}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)\right)}_{\mbox{Derivative of log density}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
Proceeding to differentiate the above quantity with respect to $x_{2}$, $x_{3}$, and $x_{4}$, provides from the product rule a sum of terms which are of the form above, but with the derivative of the log density replaced by the product of higher order derivatives of the log density.
@@
$\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)=\sum_{t=1}^{T}-\frac{1}{2\sigma_{w}^{2}}\left\|X_{t+1}-f(X_{t},\pi^{\theta_{\bm{s}}};\phi_{\bm{m}})\right\|^{2}+T\log\frac{1}{\sqrt{2\pi\sigma_{w}^{2d_{\mathsf{X}}}}}.$
By the differentiability of $f$ with respect to $\phi$ and $u$ from Assumption 2.1 and the differentiability of $\pi^{\theta}$ with respect to $\theta$ from Assumption 2.2, we have that the log density above is four times differentiable in $x_{1},\dots x_{4}$ for any values of $\Delta_{i}^{\theta}$ and $\Delta_{j}^{\phi}$.
@@
$\frac{d}{dm_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{\phi}f(X_{t},\pi^{\theta;\phi})\Delta_{i}^{\phi}$
$\frac{d}{ds_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{u}f(X_{t},\pi^{\theta};\phi)D_{\theta}\pi^{\theta}\Delta_{i}^{\theta}$
@@
$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)$
@@
×$\times\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{4}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})$
$\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
×$\times\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})},$
@@
$\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
$=\sqrt{\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]}\leq L_{\mathsf{cost}}.$
@@
$\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}\leq\mathsf{poly}(T,d_{\mathsf{X}}).$
Therefore our derivative is bounded as
$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1},L_{\mathsf{cost}},d_{\mathsf{X}}\right).$
Given Lemma B.1, the proof of Lemma 2.1 follows from an argument identical to that of Proposition 6 in Wagenmaker et al. (2023), with Lemma B.1 invoked in place of their Lemma D.1. We may also use Lemma B.1 to show the following two results from Wagenmaker et al. (2023).
@@
Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
The proofs of Footnote 3, Lemma B.2, and Lemma B.3 from the same arguments as the proofs of Lemma D.2, Lemma D.3 and Lemma D.4 of Wagenmaker et al. (2023) with Lemma B.1 taking place of their Lemma D.1, and the conclusion of Lemma 2.1 taking place of their Assumption 13.
Appendix C Modifications to Dynamic Optimal Experiment Design
C.1 Dynamic Optimal Experiment Design
We consider an experiment design procedure, Algorithm 2, which is a modified version of the algorithm from Wagenmaker et al. (2023). The goal of this algorithm is to collect a dataset consisting of $N$ episodes playing exploration policies belonging to $\Pi_{\exp}$ whose Gram matrix $\Lambda$ (15) which approximately minimizes the objective $\Phi(\Lambda)$. It takes as input the objective $\Phi$, an episode budget $N$, an initial exploration policy $\pi^{0}$, a parameter estimate $\hat{\phi}$, and a confidence set for the parameters $\mathcal{B}$. The algorithm plays a sequence of $N^{1/3}$ epochs, each with $N^{2/3}$ episodes. For the first epoch, it plays the initial policy to set an initial value of the Gram matrix $\Lambda_{0}$. Starting from this value, the algorithm performs a variant of gradient descent on the objective $\Phi$, where each epoch is a gradient step. At each iteration, it computes the gradient $\Xi_{n}$ of this objective at the current iterate $\Lambda_{n}$. Rather than taking a step in the direction $\Xi_{n}$, it takes a step in the descent direction constrained to the set of Gram matrices achievable under the dynamics evolve according to (1) when the input is chosen according to some exploration policy $\pi\in\Pi_{\exp}$. It does so by estimating the solution to
$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right],$
where $c^{n}$ is defined in terms of the gradient $\Xi_{n}$ as
$c^{n}(x,u)=\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{n}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}.$
In the above optimization problem, the expectation is taken with respect to the true parameter, which is unknown. However, the confidence ball $\mathcal{B}$ centered at the estimate $\hat{\phi}$ can be used to approximate the solution. In particular, the algorithm applies optimism in the face of uncertainty (Abbasi-Yadkori et al., 2011; Kakade et al., 2020) to select the exploration policy which solves the problem
$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\min_{\phi\in\mathcal{B}}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right].$
Doing so allows us to achieve high probability bounds on the objective value obtained by the dynamic optimal experiment design algorithm (20). The formal version of this is modified from Theorem 3 of Wagenmaker et al. (2023), and is presented in Theorem C.1.
[…]
1:Input: Objective $\Phi$, episodes $N$, initial exploration policy $\pi^{0}$, parameter estimate $\hat{\phi}$, parameter uncertainty set $\mathcal{B}$, smoothness bound $B_{\Phi}$
2:Set $\tilde{N}\leftarrow\lfloor N^{2/3}\rfloor$, $\tilde{K}\leftarrow\lfloor N^{1/3}\rfloor-1$, $\gamma_{n}\leftarrow\frac{1}{n+1}$.
3:Play $\pi^{0}$ for $\tilde{N}$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$,
4:Store $\mathcal{D}_{0}\leftarrow\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,k=1}^{T+1,\tilde{N}}$
5:Form the Gram matrix $\Delta_{0}=\frac{1}{\tilde{N}}\sum_{{t,n=1}}^{T,\tilde{N}}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})$
6:for $k=1,\dots,\tilde{K}$ do
7:     Define $\Xi_{k}\leftarrow\nabla_{\Lambda}\Phi(\Lambda)|_{\Lambda=\Lambda_{k-1}}$ and $c^{k}(x,u)\!\leftarrow\!\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{k}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}$
8:     Set $\pi^{k}\leftarrow\underset{\tilde{\pi}\in\Pi_{\exp}}{\operatorname*{argmin}}\,\underset{\phi\in\mathcal{B}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{k}(X_{t},U_{t})\right]$
9:     Run $\pi^{k}$ for $\tilde{N}$ episodes, collecting $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T,\tilde{N}}$, and update the dataset and Gram matrix as
$\Lambda_{k}$ $\leftarrow(1-\gamma_{k})\Lambda_{k-1}+\frac{\gamma_{k}}{\tilde{N}}\sum_{t,n=1}^{T,\tilde{N}}Df(x,u;\hat{\phi})^{\top}\left(\Xi_{k}\right)Df(x,u;\hat{\phi}),\quad\mathcal{D}_{k}\leftarrow\mathcal{D}_{k-1}\cup\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$
10:Return: $\Lambda_{\tilde{K}}$, $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}$
C.2 Proof of correctness for $\texttt{DOED}_{+}$ Algorithm
The objective of Algorithm 2 is to collect data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{n,t=1}^{N,T}$ from the system to optimize the some objective $\Phi(\Lambda)$, where
$\Lambda=\sum_{n,t=1}^{N,T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}).$
In light of this, we define the set of all possible Gram matrices corresponding to the parameter parameter $\hat{\phi}$:
$\Omega(\phi)$ =$=\bigg{\{}\sum_{t=1}^{T}Df(x_{t},u_{t};\phi)^{\top}Df(x_{t},u_{t};\phi):x_{t}\in\mathbb{R}^{d_{\mathsf{X}}},u_{t}\in\mathbb{R}^{d_{\mathsf{U}}},\forall h\in[H]\bigg{\}}.$
The correctness of the algorithm can be guaranteed under the following regularity assumption on the objective.
Assumption C.1.
[…]
$\left\|\nabla\Phi(\Lambda)-\nabla\Phi(\Lambda^{\prime})\right\|_{\star}\leq\kappa\left\|\Lambda-\Lambda^{\prime}\right\|\,\,\,\,\forall\Lambda,\Lambda^{\prime}\in\Omega(\hat{\phi}).$
Furthermore, suppose there exists some $B_{\Phi}<\infty$ such that for all $\Lambda\in\Omega(\hat{\phi}),x\in\mathbb{R}^{d_{\mathsf{X}}},u\in\mathbb{R}^{d_{\mathsf{U}}}$,
$\left\|Df(x,u,\hat{\phi})\nabla\Phi(\Lambda)Df(x,u,\hat{\phi})^{\top}\right\|\leq B_{\Phi}.$
This leads us to the following result, characterizing the optimality of the data collected via the dynamic optimal experiment design procedure.
Theorem C.1 (Modification of Theorem 3 of Wagenmaker et al. (2023)).
Suppose Assumptions 2.1 and C.1 hold. Additionally, suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$. Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfies
$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$\leq\frac{2\kappa T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}\!+\!B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}\!+\!\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right).$
To prove the above result, we first present an extension of the simulation lemma from Kakade et al. (2020) to the setting of smooth nonlinear systems.
Lemma C.1 (Self-Bounding, Simulation Lemma: Modified from Kakade et al. (2020)).
Suppose Assumption 2.1 holds. For any policy $\pi$, model parameters $\phi$ and non-negative cost $c$ satisfying $c<1$, it holds that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi-\phi^{\star}\right\|.$

### v2

@@ A.2 Improved Rates via the Delta Method
$\tau_{\mathsf{cov\,conc}}(\delta)=\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta))}{\beta^{2}}.$
If $K\geq\tau_{\mathsf{cov\,conc}}(\delta/4)$, we may invoke Lemma A.9 to show that with probability at least $41-\delta/$ 4,
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta.$
By the fact that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{4}$, we have that $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$. Given this condition and the burn-in conditions $K\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/8),\tau_{\mathsf{small\,error}}(\delta/4),\tau_{\mathsf{h.o.t.}}(\delta/4,\beta)\right\}\mathclose{}$, we satisfy the conditions of Lemma A.5. Therefore, conditioned on the event of covariance concentration, we have that with probability at least $41-\delta/$ 4,
@@
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$. Then by the triangle inequality, we have that under the success events of covariance concentration and the delta method,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq\left(\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}+\beta\left\|H\right\|^{1/2}\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|\right)^{2}.$ (28)
We have reduced the parameter recovery errort to the sum of weighted norms of a self-normalized martingale. Calling on self-normalized martingale machinary Abbasi-Yadkori et al. (2011). In particular, invoking Lemma A.8 with failure probability $\delta/4$ and union bounding provides that with probability at least $21-\delta/$ 2, the following bounds hold:
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}\leq\sqrt{2\sigma_{w}^{2}(1+\xi)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}(1+\xi)\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{4}{\delta}},$
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|\leq\sqrt{2\sigma_{w}^{2}(1+\xi)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1})+4\sigma_{w}^{2}(1+\xi)\left\|(K\Sigma^{\pi})^{-1}\right\|\log\frac{4}{\delta}},$
where $\xi=\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}$ Substituting these bounds into (28), we find that by union bounding over the success events, we have that with probability at least $1-\delta$,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 2\sigma_{w}^{2}(1+\xi+4\beta d_{\phi})\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}(1+\xi+4\beta d_{\phi})\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{4}{\delta}.$
@@ Appendix B Results for smooth nonlinear systems
In particular, we first define $p_{w}$ to denote the density of the noise, $W_{t}$. Let $p_{\phi,\theta}(\cdot)$ denote the density over the states $X_{2:T+1}$ induced by playing controller $\pi^{\theta}$ on the system (1) with parameter $\phi$. Observe that
$p_{\theta,\phi}(X_{2:T+1})=\prod_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi^{\theta}_{t}(X_{t});\phi^{\star})).$ (29)
@@
[…]
with $X_{1}$ fixed as 00.
@@
[…]
Proceeding to differentiate the above quantity with respect to $x_{2}$, $x_{3}$, and $x_{4}$, provides from the product rule a sum of terms which are of the form above, but with the derivative of the log density and the cost gradient replaced by the product of higher order derivatives of the log density.
@@
$\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)=\sum_{t=1}^{T}-\frac{1}{2\sigma_{w}^{2}}\left\|X_{t+1}-f(X_{t},\pi^{\theta_{\bm{s}}};\phi_{\bm{m}})\right\|^{2}+T\log\frac{1}{\sqrt{2\pi\sigma_{w}^{2d_{\mathsf{X}}}}}.$
By the differentiability of $f$ with respect to $\phi$ and $u$ from Assumption 2.1 and the differentiability of $\pi^{\theta}$ with respect to $\theta$ from Assumption 2.2, we have that the log density above is four times differentiable in $x_{1},\dots x_{4}$ for any values of $\Delta_{i}^{\theta}$ and $\Delta_{j}^{\phi}$. Similarly, by differentiability of the stage costs from Assumption 2.3 and the policy from Assumption 2.2, the cost gradient is three times differentiable in $x_{1},\dots,x_{4}$.
@@
$\frac{d}{dm_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{\phi}f(X_{t},\pi^{\theta;\phi})\Delta_{i}^{\phi}$
$\frac{d}{ds_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{u}f(X_{t},\pi^{\theta};\phi)D_{\theta}\pi^{\theta}\Delta_{i}^{\theta}$
@@
$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{2:T+1})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)$
@@
×$\times\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{4}\left(\sum_{t=1}^{T}c_{t}(X_{t},\pi_{\theta}(X_{t})+c_{T+1}(X_{T+1})\right)p_{\theta,\phi}(X_{1:T+1})dX_{2:T+1}$
[…]
@@
[…]
We may similarly bound the cost gradient by a polynomial in $L_{\mathsf{cost}}$ and $L_{\theta}$. Therefore our derivative is bounded as
$\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(T,L_{f},L_{\theta},\sigma_{w}^{-1},L_{\mathsf{cost}},d_{\mathsf{X}}\right).$
Given Lemma B.1, Assumption 2.4 may be shown to hold if $\nabla_{\theta}^{2}J(\pi^{\theta},\phi^{\star})\succ 0$ using an argument identical to that of Proposition 6 in Wagenmaker et al. (2023), with Lemma B.1 invoked in place of their Lemma D.1. We may also use Lemma B.1 to show the following two results from Wagenmaker et al. (2023).
@@
Under Assumptions 2.1, 2.2, 2.3 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
Under Assumptions 2.1, 2.2,2.3 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
The proofs of Lemma 3.1, Lemma B.2, and Lemma B.3 from the same arguments as the proofs of Lemma D.2, Lemma D.3 and Lemma D.4 of Wagenmaker et al. (2023) with Lemma B.1 taking place of their Lemma D.1.
Finally, we present a result that demonstrates that under our smoothness assumptions, the Fisher information is differentiable with respect to the dynamics parameter, and that the derivative is bounded by a quantity which is polynomial in relevant system parameters and the length of the horizon.
Lemma B.4.
For any exploration policy $\pi\in\Pi_{\exp}$, any $\phi\in\mathbb{R}^{d_{\phi}}$, and any unit vector $v\in\mathbb{R}^{d_{\phi}}$ we have that $v^{\top}\mathbf{FI}^{\pi}(\phi)v$ is differentiable with respect to $\phi$, and that
$\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v\right\|\leq 3\frac{T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}.$